A human bone is subjected to a compressive force of 5.0×10^5 N. The bone is 25 cm long and has approximate cross sectional area of 4.0 cm^2. If the ultimate compressive strength of the bone is 1.70×10^8 N/m^2, will the bone be compressed or it will break under this force ?

Grade:11

1 Answers

Arun
6 years ago
Stress = F/A
= 5 * 10^5/ 4 * 10^-4
 = 12.5 * 10^8
 
Since this stress exceeds the ultimate compressive stress of the bone,
 
Hence bone will break.
 
Regards
Arun (askIITians forum expert)